OpenStudy (randomecorner):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

OpenStudy (dan815):

what will the expansion of 5(9-3) look like

OpenStudy (dan815):

remember the distributive law of multiplication

OpenStudy (randomecorner):

30

OpenStudy (dan815):

yea okay but i mean just writing it out in general

OpenStudy (dan815):

like the middle step

OpenStudy (dan815):

5(9-3)=5*9 - 5*3

OpenStudy (randomecorner):

cuz of 9-3=6 times 5=30

OpenStudy (dan815):

okay dont simplify the inside now, just write out distributive law

OpenStudy (dan815):

5(9 − 3) = w • 9 – 5 • 3 5(9 − 3) = 5 * 9 - 5 * 3 <------- from distributive law of multiplication can you see what w has to be now

OpenStudy (dan815):

you can also do it this way if u want you already know it has to be 30

OpenStudy (randomecorner):

my math bokis horrible at explaining

OpenStudy (dan815):

30=w • 9 – 5 • 3 so what does w have to be to satisfy this equation

OpenStudy (dan815):

you know the period there is for multiplication right

OpenStudy (randomecorner):

by its self?

OpenStudy (dan815):

30=w • 9 – 5 • 3 hmm okay

OpenStudy (randomecorner):

and yes i do know that

OpenStudy (dan815):

look try plugging in numbers to get a feel for it

OpenStudy (dan815):

lets say u make w=2

OpenStudy (dan815):

is 2*9-5*3 = 30?

OpenStudy (randomecorner):

no

OpenStudy (dan815):

right so w is not 2

OpenStudy (dan815):

try a bigger number for w

OpenStudy (dan815):

lets say w=4

OpenStudy (dan815):

is this right 4*9-5*3=30?

OpenStudy (dan815):

are getting closer atleast

OpenStudy (randomecorner):

yes

OpenStudy (dan815):

try w=5 now what happens?

OpenStudy (randomecorner):

or i miscalculated it

OpenStudy (randomecorner):

no its still wrng

OpenStudy (randomecorner):

w=6

OpenStudy (randomecorner):

is and itwrong also

OpenStudy (randomecorner):

i hate math

OpenStudy (randomecorner):

5(6)= w(4)(3) 30=12w 30/12=w but 30/12 cant get enterd in the bx so we divide?

OpenStudy (randomecorner):

wich is 2.5

OpenStudy (randomecorner):

but we can break it down more to 1 by dividing them right?